23. Suppose 10 out of 25 students in your class have dogs for pets.
a. What is the theoretical probability that a student selected at random from your class has a dog?
b. Using your result from part (a), predict how many dog-owning students you can expect among the 200 students in your grade. Show your work.

Answer by chessace(471)

a. p=10/25 = 0.4
b. expect p*total = 0.4 * 200 = 80